# Automatic Speech Recognition using Advanced Deep Learning Approaches: A survey

### v1

@@ 3.1 Acoustic and language models
$P(w_{n}|w_{n-1},w_{n-2},\ldots,w_{n-(n-1)})=\frac{C(w_{n-(n-1)},\ldots,w_{n})}{C(w_{n-(n-1)},\ldots,w_{n-1})}$ (2)
@@ 3.2 Evaluation criteria in ASR
To assess the effectiveness and suitability of ASR techniques, researchers have employed diverse methods. Some of these encompass well-established DL metrics, including accuracy, F1-score, recall (sensitivity or true positive rate (TPR)), precision (also known as positive predictive value), and specificity (commonly referred to as true negative rate (TNR)) [41]. These metrics serve as crucial evaluation criteria for experimental outcomes, as evidenced in studies such as [42, 43, 44]. Additionally, there are ASR-specific metrics, which are detailed in Table LABEL:tab:metrics.
Metric
Formula
Description
WER
$\mathrm{\frac{S+D+I}{N}=\frac{S+D+I}{H+S+D}}.$
The word error rate (WER) serves as a frequently utilized metric to assess the performance of Automatic Speech Recognition (ASR). It is computed by determining the ratio of incorrectly recognized words to the overall number of processed words [45, 17, 46]. In the given context, $\mathrm{I,D,S,H}$, and $\mathrm{N}$ denote the quantities of insertions, deletions, substitutions, hits, and input words, respectively. Instead of WER, the character error rate (CER) has been employed, while adhering to the same evaluation principle.
PESQ and MOS-LQO
$\mathrm{MOS-LQO}=0.999+\frac{4.999-0.999}{1+e^{-1.4945.PESQ+4.6607}}$
PESQ serves as an objective technique for evaluating the perceived quality of speech [47]. The assessment involves assigning numerical scores within the range of -0.5 to 4.5. Additionally, a correlation can be established between MOS and PESQ scores, giving rise to a novel evaluation metric termed the mean opinion score-listening quality objective (MOS-LQO), also identified as PESQ Rec.862.1. [22]
RTF
RTF=Total Processing TimeTotal $\mathrm{RTF=\frac{\text{Total Processing Time}}{\text{Total Duration}}}$
[…]
RTF serves as a standard metric to assess the processing time cost of an ASR system. It represents the average processing time required for one second of speech
RER
$\frac{(E_{\text{baseline}}-E_{\text{proposed}})}{E_{\text{baseline}}}\times 100\%$
The relative error rate (RER) expresses the percentage error rate achieved by the proposed DL model compared to the baseline. $E_{\text{baseline}}$ is the error rate of the baseline model. $E_{\text{proposed}}$ is the error rate of the proposed model or method.
D
$\frac{\sum_{i=1}^{n}\Big{(}1-\frac{\sum_{j=1}^{n}a_{ij}\cdot|i-j|}{\max(|i-1|,|i-2|,\ldots,|i-n|)}\Big{)}}{n}\hskip 28.45274pt$
Diagonal centrality of an attention matrix (D) is defined as the mean value across the centrality of all its rows. where $j$ represents the index of each column, $n$ signifies the length of the input sequence, $a_{ij}$ denotes the attention weight between the $i$-th and $j$-th elements of the input sequence, and $|i-j|$ signifies the distance between the $i$-th and $j$-th elements of the input sequence [48].
@@ 3.3 ASR datasets
. Dataset Used by Default ASR task Characteristics LibriSpeech [49, 50, 51] Train and assess systems for recognizing speech. The collection consists of 1000 hours of speech recorded at a 16 kHz sampling rate, sourced from audiobooks included in the LibriVox project. DCASE [52, 53] Identifying acoustic environments and detecting sound occurrences. Comprise 8 coarse-level and 23 fine-level urban sound categories, collected in New York City in 2020 using 50 acoustic sensors. WSJ [48] Acoustic scene and sound event corpus Comprises an extensive 81 hours of meticulously curated read speech training data. SWBD [48] Conversational telephone speech corpus Is a comprehensive collection, boasting a substantial 260 hours of conversational telephone speech training data. AISHELL [54, 55, 56] Chinese Mandarin speech corpus 400 participants from diverse Chinese accent regions recorded in a quiet indoor space using high-fidelity microphones, later downsampled to 16kHz. CHIME3 [57] SR for distant microphone in real-world settings. Includes around 342 hours of English speech with noisy transcripts and 50 hours of noisy environment recordings. Google-SC [57] Speech commands with a restricted range of words. The dataset contains 105,829 one-second utterances of 35 words categorized by frequency. Each utterance is stored as a one-second WAVE format file with 16-bit single-channel at 16KHz rate. It involves 2,618 speakers AURORA4 [57] Compare front-ends for large vocabulary recognition performance. Aurora-4 is a speech recognition dataset derived from the WSJ corpus, offering four conditions (Clean, channel, noisy, channel+noisy) with two microphone types and six noise types, totaling 4,620 utterances per set. Car-env [57] Vehicle environment sound Is a dataset from Korea that spans 100 hours of recordings in a vehicle. It comprises brief commands, with an average of 1.6 words per command. HKUST [50, 56] Classify Mandarin speech into standard and accented types. Comprises roughly 149 hours of telephone conversations in Mandarin. AudioSet [51] Audio event recognition Includes 1,789,621 segments of 10 seconds each (equivalent to 4,971 hours). It consists of at least 100 instances clustered into 632 audio classes, with only 485 audio event categories clearly identified. AWIC-19 [58] Arabic words recognition It comprises 770 recordings featuring isolated Arabic words. TED2 [59] English corpus for ASR The dataset was first made available in May 2012, for training, it comprising 118 hours, 4 minutes, and 48 seconds of training data from 666 speakers, containing approximately 1.7 million words.
@@ 4 Advanced ASR methods and applications
Traditional statistical LMs, such as backoff n-gram LMs, have been widely used due to their simplicity and reliability. However, bidirectional encoder representations from transformers (BERT), which utilize attention models, have shown better contextual understanding compared to single-direction LMs, as demonstrated in the work of Devlin et al. [60].
In terms of AM, deep learning-based models like the deep neural network-hidden Markov model (DNN-HMM) and the connectionist temporal classification (CTC) have made significant advancements. DNN-HMM models have been extensively studied in ASR research, while CTC is an end-to-end training method that does not require pre-alignment and only needs input and output sequences. The S2S model has also been successful in solving ASR tasks without using an LM or pronunciation dictionary, as described in Chiu et al. [61].
@@ 4.1 Transformer-based ASR
The Transformer stands as a prominent deep learning model extensively employed across diverse domains, including NLP, computer vision (CV), and speech processing. Originally conceived for machine translation as a S2S model, it has evolved to find applications in various fields. The Transformer heavily relies on the self-attention mechanism, enabling it to capture extensive dependencies in input sequences. The standard Transformer model incorporates the query–key–value (QKV) attention mechanism. In this setup, given matrix representations of queries $\mathbf{Q}\in\mathbb{R}^{N\times D_{k}}$, keys $\mathbf{K}\in\mathbb{R}^{M\times D_{k}}$, and values $\mathbf{V}\in\mathbb{R}^{M\times D_{v}}$, the scaled dot-product attention is defined as:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{D_{k}}}\right)\mathbf{V}$
@@
  • Self-attention: In the Transformer encoder, queries $\mathbf{Q}$, keys $\mathbf{K}$, and values $\mathbf{V}$ all equal the outputs of the previous layer, denoted as $\mathbf{X}$ in Eq. (2).
@@
Numerous studies in the ASR field have introduced transformer-based approaches, encompassing both the acoustic and language domains. In the subsequent subsections, we delve into a comprehensive review and detailed analysis of several cutting-edge techniques within each of these categories.
@@
Ref.
Based on
Speech recognition task
Transformer
AM/LM
Result with metric
[55]
CNN
Solve the problem of code-switching
Multi-head attention
LM
RER= 10.2%
[56]
VGGnet
Compress ASR parameters and speeds up the inference time
Low-rank multi-head attention
AM
CER= 13.09%
[57]
DNN-HMM
Improve ASR
Attention
AM
RER= 4.7$\downarrow$
[62]
Emformer
Large scale ASR
Attention
AM
RERR= 26%
[63]
TRUNet
Sound source separation
TNet
AM
PESQ= 0.22$\uparrow$
[64]
MHSA
Improve speech/ASR
[…]
AM
PESQ= 0.96$\uparrow$
[58]
HMM
Improve ASR
Acoustic Encoder
AM
Acc= 96%
[65]
RNN-T
Acoustic re-scoring scenario
Transformer- Transducer
AM
Acc= 8$\uparrow$
[66]
CTC
ASR, ST, Acoustic ED
All-in-one
AM
WER=0.3$\uparrow$
[67]
CNN
Speech recognition with low latency, reduced frame rate, and streamability.
Transformer-Transducer
AM
WER= 3.6%
[68]
CTC alignment
Retrieve the acoustic embedding at the token level for better ASR
Attention
AM
@@
[69]
RNN-LSTM
Improve the efficiency of end-to-end ASR
Attention
LM
CER=1.98%
[59]
CTC Alignment
Enhance the performance of end-to-end ASR
Autoregressive Transformer
AM
RTF= 0.0134
@@ 4.1.1 Acoustic domain
The study [57] reveals the Transformer model’s increased susceptibility to input sparsity compared to the convolutional neural network (CNN). The authors analyze the performance decline, attributing it to the Transformer’s structural characteristics. Additionally, they introduce a novel regularization method to enhance the Transformer’s resilience to input sparsity. This method directly regulates attention weights through silence label information in forced-alignment, offering the advantage of not requiring extra module training and excessive computation.
@@
The proposed masked autoencoding audio spectrogram Transformer (MAE-AST) operates solely on unmasked tokens [51], utilizing a large encoder. It concatenates mask tokens with encoder output embeddings, feeding them into a shallow decoder. Fine-tuning for downstream tasks involves using only the encoder, eliminating the decoder’s reconstruction layers. MAE-AST represents a significant improvement over the self-supervised audio spectrogram transformer (SSAST) model for speech and audio classification. Addressing the high masking ratio issue, the method achieves a 3$\times$ speedup and 2$\times$ memory usage reduction. During downstream tasks, the approach consistently outperforms SSAST. To identify varities of sounds types, Bai et al. [52] introduce SE-Trans, a cross-task model for environmental sound recognition, encompassing acoustic scene classification, urban sound tagging, and anomalous sound detection. Utilizing attention mechanisms and Transformer encoder modules, SE-Trans learns channel-wise relationships and temporal dependencies in acoustic features. The model incorporates FMix data augmentation, involving the creation of a binary mask from a randomly sampled complex matrix with a low-pass filter. SE-Trans achieves outstanding performance in ESR tasks, proven through evaluations on DCASE challenge databases, underscoring its robustness and versatility in environmental sound recognition.
Automated audio captioning (AAC) involves generating textual descriptions for audio recordings, covering sound events, acoustic scenes, and event relationships. Current AAC systems typically employ an encoder-decoder architecture, with the decoder crafting captions based on extracted audio features. Chen et al. in their paper [53] introduces a novel approach that enhances caption generation by leveraging multi-level information extracted from the audio clip. The proposed method consists of a CNN encoder with multi-level feature extraction (channel attention, spatial attention), A module specialized in predicting keywords to generate guidance information at the word level and Transformer decoder. Figure 6 depicts the overall architecture incorporating the three mentioned modules. Results demonstrate significant improvements in various metrics, achieving SOTA performance during the cross-entropy training stage.
@@
TRUNet proposed in [63] represents an innovative approach to end-to-end multi-channel reverberant sound source separation. The model incorporates a recurrent-U network that directly estimates multi-channel filters from input spectra, enabling the exploitation of spatial and spectro-temporal diversity in sound source separation. In Figure 7, the block diagram illustrates the architecture of TRUNet’s transformer. The transformer network (TNet) encompasses three variations: (i) TNet–Cat, which concatenates multi-channel spectra, treating them as a single input. This approach allows for the direct utilization of spatial information between channels. (ii) TNet–RealImag, utilizing two separate transformer stacks for real and imaginary parts, respectively. Queries and keys are computed from the multi-channel spectra. Despite this, the method may not fully exploit spatial information, such as phase differences, directly. (iii) TNet–MagPhase, analogous to TNet–RealImag, but employing spectral magnitude and spectral phase instead of real and imaginary parts. This variation proves superior in extracting spatial information from complex-valued spectra, resulting in maximum enhanced performance in sound source separation when employing TRUNet-MagPhase architecture.

### v2

@@ 2.1 Acoustic and language models
$\scriptsize{P(w_{n}|w_{n-1},w_{n-2},\ldots,w_{n-(n-1)})=\frac{C(w_{n-(n-1)},\ldots,w_{n})}{C(w_{n-(n-1)},\ldots,w_{n-1})}}$ (1)
@@ 2.2 Evaluation criteria in ASR
To assess the effectiveness and suitability of ASR techniques, researchers have employed diverse methods. Some of these encompass well-established DL metrics, including accuracy, F1-score, recall (sensitivity or true positive rate (TPR)), precision (also known as positive predictive value), and specificity (commonly referred to as true negative rate (TNR)) [41]. These metrics serve as crucial evaluation criteria for experimental outcomes, as evidenced in studies such as [42, 43, 44]. Additionally, there are ASR-specific metrics, which are detailed in Table 2.
Table 2: An overview of the metrics employed for evaluating ASR methods.
Metric Formula Description
WER $\mathrm{\frac{S+D+I}{N}=\frac{S+D+I}{H+S+D}}.$ The word error rate (WER) serves as a frequently utilized metric to assess the performance of Automatic Speech Recognition (ASR). It is computed by determining the ratio of incorrectly recognized words to the overall number of processed words [45, 17, 46]. In the given context, $\mathrm{I,D,S,H}$, and $\mathrm{N}$ denote the quantities of insertions, deletions, substitutions, hits, and input words, respectively. Instead of WER, the character error rate (CER) has been employed, while adhering to the same evaluation principle.
PESQ and MOS-LQO $\mathrm{MOS-LQO}=0.999+\frac{4.999-0.999}{1+e^{-1.4945.PESQ+4.6607}}$ PESQ serves as an objective technique for evaluating the perceived quality of speech [47]. The assessment involves assigning numerical scores within the range of -0.5 to 4.5. Additionally, a correlation can be established between MOS and perceptual evaluation of speech quality (PESQ) scores, giving rise to a novel evaluation metric termed the mean opinion score-listening quality objective (MOS-LQO), also identified as PESQ Rec.862.1. [22]
RTF RTF=Total Processing TimeTotal $\mathrm{RTF=\frac{\text{Total Processing Time}}{\text{Total Duration}}}$Ac RTF serves as a standard metric to assess the processing time cost of an ASR system. It represents the average processing time required for one second of speech
RER $\frac{(E_{\text{baseline}}-E_{\text{proposed}})}{E_{\text{baseline}}}\times 100\%$ The relative error rate (RER) expresses the percentage error rate achieved by the proposed DL model compared to the baseline. $E_{\text{baseline}}$ is the error rate of the baseline model. $E_{\text{proposed}}$ is the error rate of the proposed model or method.
D $\frac{\sum_{i=1}^{n}\Big{(}1-\frac{\sum_{j=1}^{n}a_{ij}\cdot|i-j|}{\max(|i-1|,|i-2|,\ldots,|i-n|)}\Big{)}}{n}\hskip 28.45274pt$ Diagonal centrality of an attention matrix (D) is defined as the mean value across the centrality of all its rows. where $j$ represents the index of each column, $n$ signifies the length of the input sequence, $a_{ij}$ denotes the attention weight between the $i$-th and $j$-th elements of the input sequence, and $|i-j|$ signifies the distance between the $i$-th and $j$-th elements of the input sequence [48].
@@ 2.3 ASR datasets
. Dataset Used by Default ASR task Characteristics LibriSpeech [49, 50, 51] Train and assess systems for recognizing speech. The collection consists of 1000 hours of speech recorded at a 16 kHz sampling rate, sourced from audiobooks included in the LibriVox project. DCASE [52, 53] Identifying acoustic environments and detecting sound occurrences. Comprise 8 coarse-level and 23 fine-level urban sound categories, collected in New York City in 2020 using 50 acoustic sensors. WSJ [48] Acoustic scene and sound event corpus Comprises an extensive 81 hours of meticulously curated read speech training data. SWBD [48] Conversational telephone speech corpus Is a comprehensive collection, boasting a substantial 260 hours of conversational telephone speech training data. AISHELL [54, 55, 56] Chinese Mandarin speech corpus 400 participants from diverse Chinese accent regions recorded in a quiet indoor space using high-fidelity microphones, later downsampled to 16kHz. CHIME3 [57] SR for distant microphone in real-world settings. Includes around 342 hours of English speech with noisy transcripts and 50 hours of noisy environment recordings. Google-SC [57] Speech commands with a restricted range of words. The dataset contains 105,829 one-second utterances of 35 words categorized by frequency. Each utterance is stored as a one-second WAVE format file with 16-bit single-channel at 16KHz rate. It involves 2,618 speakers Aurora-4 [57] Compare front-ends for large vocabulary recognition performance. Aurora-4 is a speech recognition dataset derived from the WSJ corpus, offering four conditions (Clean, channel, noisy, channel+noisy) with two microphone types and six noise types, totaling 4,620 utterances per set. Car-env [57] Vehicle environment sound Is a dataset from Korea that spans 100 hours of recordings in a vehicle. It comprises brief commands, with an average of 1.6 words per command. HKUST [50, 56] Classify Mandarin speech into standard and accented types. Comprises roughly 149 hours of telephone conversations in Mandarin. AudioSet [51] Audio event recognition Includes 1,789,621 segments of 10 seconds each (equivalent to 4,971 hours). It consists of at least 100 instances clustered into 632 audio classes, with only 485 audio event categories clearly identified. AWIC-19 [58] Arabic words recognition It comprises 770 recordings featuring isolated Arabic words. TED2 [59] English corpus for ASR The dataset was first made available in May 2012, for training, it comprising 118 hours, 4 minutes, and 48 seconds of training data from 666 speakers, containing approximately 1.7 million words.
@@ 3 Advanced ASR methods and applications
Traditional statistical LMs, such as backoff n-gram LMs, have been widely used due to their simplicity and reliability. However, bidirectional encoder representations from transformers (BERT) [60], which utilize attention models, have shown better contextual understanding compared to single-direction LMs, as demonstrated in the work of Devlin et al. [61].
In terms of AM, deep learning-based models like the deep neural network-hidden Markov model (DNN-HMM) and the connectionist temporal classification (CTC) have made significant advancements [60]. DNN-HMM models have been extensively studied in ASR research, while CTC is an end-to-end training method that does not require pre-alignment and only needs input and output sequences. The S2S model has also been successful in solving ASR tasks without using an LM or pronunciation dictionary, as described in Chiu et al. [62].
@@
The figures in this section are designed to fulfill two key objectives: firstly, to visually explain the principles behind advanced DL techniques, namely Transformers, DTL, FL, and DRL, as depicted in Figures 5, 8, 9, and 10, respectively; and secondly, to present practical examples of implementing widely recognized techniques. These include a convolutional neural network (CNN)-based Transformer illustrated in Figure 6, a particularly effective ASR application of source separation using a Transformer, highlighted in Figure 7, and a S2S-based ASR model depicted in Figure 11. This dual presentation provides young researchers and developers with valuable, practical insights on how to implement these advanced DL techniques in their projects. By doing so, it bridges the gap between theory and practice, encouraging the application and extension of these examples in new and innovative ways.
@@ 3.1 Transformer-based ASR
The Transformer stands as a prominent deep learning model extensively employed across diverse domains, including NLP, computer vision (CV), and speech processing. Originally conceived for machine translation as a S2S model, it has evolved to find applications in various fields. The Transformer heavily relies on the self-attention mechanism, enabling it to capture extensive dependencies in input sequences. The standard Transformer model incorporates the query–key–value (QKV) attention mechanism. In this setup, given matrix representations of queries $\mathbf{Q}\in\mathbb{R}^{N\times D_{k}}$, keys $\mathbf{K}\in\mathbb{R}^{M\times D_{k}}$, and values $\mathbf{V}\in\mathbb{R}^{M\times D_{v}}$, the scaled dot-product attention is defined as formulated in Equation 2:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{D_{k}}}\right)\mathbf{V}$ (2)
@@
  • Self-attention: In the Transformer encoder, the queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) are all equal to the outputs of the previous layer, considered as $\mathbf{X}$, or to the initial embeddings in the case of the first layer.
@@
Numerous studies in the ASR field have introduced Transformer-based approaches, encompassing both the acoustic and language domains. In the subsequent subsections, we delve into a comprehensive review and detailed analysis of several cutting-edge techniques within each of these categories. Table 4 summarises the most recent Transformer-based ASR techniques used in AM and LM domains.
@@
Ref. Based on Speech recognition task Transformer AM/LM Result with metric
[55] CNN Solve the problem of code-switching Multi-head attention LM RER= 10.2%
[56] VGGnet Compress ASR parameters and speeds up the inference time Low-rank multi-head attention AM CER= 13.09%
[57] DNN-HMM Improve ASR Attention AM RER= 4.7$\downarrow$
[63] Emformer Large scale ASR Attention AM RERR= 26%
[64] TRUNet Sound source separation TNet AM PESQ= 0.22$\uparrow$
[65] MHSA Improve speech/ASR D2Net AM PESQ= 0.96$\uparrow$
[58] HMM Improve ASR Acoustic Encoder AM Acc= 96%
[66] RNN-T Acoustic re-scoring scenario Transformer- Transducer AM Acc= 8$\uparrow$
[67] CTC ASR, ST, Acoustic ED All-in-one AM WER=0.3$\uparrow$
[68] CNN Speech recognition with low latency, reduced frame rate, and streamability. Transformer-Transducer AM WER= 3.6%
@@
[70] RNN-LSTM Improve the efficiency of end-to-end ASR Attention LM CER=1.98%
[59] CTC Alignment Enhance the performance of end-to-end ASR Autoregressive Transformer AM RTF= 0.0134
@@ 3.1.1 Acoustic domain
Figure 5: Three forms of end-to-end Transformers models: (a) attention, (b) RNN-Transducer, and (c) basic CTC [71].
The study [57] reveals the Transformer model’s increased susceptibility to input sparsity compared to the CNN. The authors analyze the performance decline, attributing it to the Transformer’s structural characteristics. Additionally, they introduce a novel regularization method to enhance the Transformer’s resilience to input sparsity. This method directly regulates attention weights , i,e. output of Figure 5 (a), through silence label information in forced-alignment, offering the advantage of not requiring extra module training and excessive computation.
@@
The proposed masked autoencoding audio spectrogram Transformer (MAE-AST) operates solely on unmasked tokens [51], utilizing a large encoder. It concatenates mask tokens with encoder output embeddings, feeding them into a shallow decoder. Fine-tuning for downstream tasks involves using only the encoder, eliminating the decoder’s reconstruction layers. MAE-AST represents a significant improvement over the self-supervised audio spectrogram transformer (SSAST) model for speech and audio classification. Addressing the high masking ratio issue, the method achieves a 3$\times$ speedup and 2$\times$ memory usage reduction. During downstream tasks, the approach consistently outperforms SSAST. To identify varities of sounds types, Bai et al. [52] introduce SE-Trans, a cross-task model for environmental sound recognition, encompassing acoustic scene classification, urban sound tagging, and anomalous sound detection. Utilizing attention mechanisms and Transformer encoder modules, SE-Trans learns channel-wise relationships and temporal dependencies in acoustic features. The model incorporates FMix data augmentation, involving the creation of a binary mask from a randomly sampled complex matrix with a low-pass filter. SE-Trans achieves outstanding performance in ASR tasks, proven through evaluations on DCASE challenge databases, underscoring its robustness and versatility in environmental sound recognition.
[…]
AAC involves generating textual descriptions for audio recordings, covering sound events, acoustic scenes, and event relationships. Current automated audio captioning (AAC) systems typically employ an encoder-decoder architecture, with the decoder crafting captions based on extracted audio features. Chen et al. in their paper [53] introduces a novel approach that enhances caption generation by leveraging multi-level information extracted from the audio clip. The proposed method consists of a CNN encoder with multi-level feature extraction (channel attention, spatial attention), A module specialized in predicting keywords to generate guidance information at the word level and Transformer decoder. Figure 6 depicts the overall architecture incorporating the three mentioned modules. Results demonstrate significant improvements in various metrics, achieving SOTA performance during the cross-entropy training stage.
@@
[…]
In recent times, dual-path networks have demonstrated effective results in many speech processing tasks such as speech separation and SE. In light of this, Wang Ke and colleagues incorporated the Transformer into the structure of dual-path networks, presenting a time-domain SE model called two-stage Transformer-based neural network (TSTNN). This model significantly enhances the performance of SE [76]. Some research findings suggest that the dot-product self-attention may not be essential for Transformer models. Similarly, the paper [65] introduces D2Net, a denoising and dereverberation network for challenging single-channel mixture speech in complex acoustic environments. D2Net incorporates a two-branch encoder (TBE) for feature extraction and fusion, along with a global-local dual-path Transformer (GLDPT) featuring local dense synthesizer attention to enhance local information perception.
@@
Shareef et al. in [58] propose a collaborative training method for acoustic encoders in Arabic DTL systems for speech-impaired children, achieving a 10% relative accuracy improvement on phoneme alignment in the output sequence. Pioneering in recognizing impaired children’s Arabic speech. Similarly in [79], collaboratively training acoustic encoders of various sizes for on-device ASR improves efficiency and reduces redundancy. Using co-distillation with an auxiliary task, collaborative training achieves up to 11% relative WER improvement on LibriSpeech corpus.
Transducer models (Figure 5 (b)), in the context of ASR, map input sequences (acoustic features) to output sequences (transcriptions). Unlike traditional S2S models, transducers can handle variable-length input and output sequences more efficiently. The study [66] explores attention masking in Transformer-Transducer-based ASR, comparing fixed masking with chunked masking in terms of accuracy and latency. The authors claim that variable masking is the viable choice in acoustic rescoring scenarios. Similarly, to adapt the Transformer for streaming DTL, the authors in [68] employ the Transducer framework for streamable alignments. Using a unidirectional Transformer with interleaved convolution layers for audio encoding, they model future context and gradually downsample input to reduce computation cost, while limiting history context length.
Moving on, the work [67] introduces an all-in-one AM based on the Transformer architecture, combined with the CTC to ensure a sequential arrangement and utilize timing details. It addresses ASR, AT, and acoustic ED simultaneously. The model demonstrates superior performance, showcasing its suitability for comprehensive acoustic scene transcription. Winata et al. [56] propose a memory-efficient Transformer architecture for end-to-end speech recognition. It significantly reduces parameters, boosting training speed by over 50% and inference time by 1.35$\times$ compared to baseline. Experiments show better generalization, lower error rates, and outperformance of existing schemes without external language or acoustic models. Growing demand for on-device ASR systems prompts interest in model compression.
3.1.2 Language domain
Self-attention models, such as Transformers, excel in speech recognition and reveal an important pattern. As upper self-attention layers are replaced with feed-forward layers, resembling convolutional long-short term deep neural network (CLDNN) architecture in [48], experiments on wall street journal (WSJ) and switchboard (SWBD) datasets show no performance drop and minor gains. The novel proposed metric of attention matrix diagonality indicates increased diagonality in lower to upper encoder self-attention layers. The authors conclude that a global view appears unnecessary for training upper encoder layers in speech recognition Transformers when lower layers capture sufficient contextual information. The study conducted by Hrinchuk et al. [49] presents a proficient postprocessing model for ASR with a Transformer-based encoder-decoder architecture, initialized with the weights of pre-trained BERT model [36]. The model effectively refines ASR output, demonstrating substantial performance gains through strategies like extensive data augmentation and pretrained weight initialization. On the LibriSpeech benchmark dataset, significant reductions in WERs are observed, particularly on noisier evaluation dataset portions, outperforming baseline models and approaching the performance of Transformer-XL neural language model re-scoring with 6-gram.